RC meeting and Run Plan, Monday November 1st

From SBS wiki
Jump to: navigation, search

Meetings Page

RC Meeting Information

RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934

Accelerator Status

Nothing new wrt accelerator, i.e. also no change of schedule (maintenance day on Wednesday, Nov. 3rd, configuration change starting on Friday Nov. 5th through the weekend, resuming operations on Monday Nov. 8th, Owl

Hall A Status and News

All shift cancelled till Friday Nov. 5th Owl; Shifts from Friday Nov. 5th owl till Sunday Nov. 7th Swing on standby; Asked Dave about when to do the target condensation, he replied that he wouldn't make firm plans before this Thursday, to see how the work in the hall has progressed (which makes sense); Since Jessie and his team were out of the hall this weekend, A few detector experts could switch their detectors on and make some tests e.g. GEMs

Subsystem Updates

  1. HCAL: took cosmics to check if TDC is fixed (yes) and check latency; latency had moved (see DAQ) and has been adjusted summary here;
  2. BBCal: HVs turned on, issues found, related to TI (see DAQ);
  3. Timing Hodoscope:
  4. Grinch: window adjusted to catch both LED signal and PMT signal;
  5. GEMS: Checked HVs stability on N2 over the weekend summary here now want to take one night of cosmics data to check the latency as well
  6. Beamline: Need to think of scheduling BCM calibration after we start running again (shall take ~1/2 shift);
  7. DAQ: TI Firmware upgrade has moved latency by 75ns
  8. Replay Analysis: discussion on Friday about mass production: held off till new version of the analyzer is ready; will need to skim out many 2D histograms (most for GEMs) which take a massive memory footprint; reviewed optics analysis; beam scalers in output tree;